OpenStudy (anonymous):

Solve using square roots x^2-81=0 1. square root of -9, square root of 9 2.-81,81 3.-9,9 4. no real number solutions

Parth (parthkohli):

snipars is love

Parth (parthkohli):

snipars is life

Parth (parthkohli):

\[x^2 = 81 = 9^2 = (-9)^2\]

Parth (parthkohli):

\[x = 9~\rm or~-9\]

OpenStudy (anonymous):

you're a god m8. thanks. ( ͡° ͜ʖ ͡°)

OpenStudy (anonymous):

that did not turn out as planned

OpenStudy (skullpatrol):

Do you understand the answer?

OpenStudy (anonymous):

I was just kerfluzzeled about the question. it's all good now.

OpenStudy (anonymous):

9. Seems as though the teacher would throw something harder in the homework though. She usually does make it a bit more challenging. though it is nice to get a break.

Parth (parthkohli):

it's all ogre now.